I am four times as old as my daughter. In 20 years time I will be as twice as old as she. How old are we now?
Answer by CharlesG2(834)

I am four times as old as my daughter. In 20 years time I will be as twice as old as she. How old are we now?
x = daughter
4x = you
x + 20 = daughter in 20 years
4x + 20 = you in 20 years
4x + 20 = 2(x + 20)
4x + 20 = 2x + 40
2x = 20
x = 10
you : 40, daughter 10
in 20 years: you: 60, daughter 30